OpenStudy (anonymous):

A floogle company's revenue is given by the function r(x)= 200- 1600/x+8-x where R is measured in thousands of dollars. Find the value of x which results in the maximum revenue. x =?

OpenStudy (campbell_st):

I assume r(x) is \[r(x) = 200 -1600/(x+8) - x\] then differentiating \[r'(x) = 1600/(x+8)^2 -1\] let r'(x) = 0 and solve for x \[1 = 1600/(x+8)^2\] \[(x + 8)^2 = 1600\] \[x + 8 = \pm \sqrt{1600}\] \[x + 8 = \pm 400\] \[x = 392, - 408\] sub each value into r(x) then maximum will be the larger r(x) after substiution